OpenStudy (anonymous):

A right angle triangle has a side x,y,z .The length of the two perpendicular sides satisfy the inequality [sqrt(x^2 - 2sqrt(2) + 9) ] + [sqrt(y^2 - 4sqrt(2) +15] <= 2sqrt(7).What is the length of hypotenuse of the triangle? A. sqrt(15) B. sqrt(10) C sqrt(7) D. sqrt(12)

OpenStudy (anonymous):

Is this the inequality \[ \sqrt{x^2 - 2\sqrt2 + 9} + \sqrt{y^2 - 4\sqrt2 +15}\leq 2\sqrt7)\ ? \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sqrt(10) i think

OpenStudy (anonymous):

can you please explain your approach

OpenStudy (anonymous):

\( 2\sqrt{7} \le \sqrt{(x-\sqrt{2})^2+7}+\sqrt{(y-2\sqrt{2})^2+7} \le2\sqrt{7} \) since \( (x-\sqrt{2})^2 \ge0 \) and \( (y-2\sqrt{2})^2 \ge0 \) so we have \( \sqrt{(x-\sqrt{2})^2+7}+\sqrt{(y-2\sqrt{2})^2+7} =2\sqrt{7} \) and equality occurs when \( x=\sqrt{2} \) and \( y=2\sqrt{2} \)

OpenStudy (anonymous):

how is it possible that \[\sqrt{(x-\sqrt{2})^{2}+7} = \sqrt{x ^{2}-2\sqrt{2}+9}\]

OpenStudy (anonymous):

instead -2sqrt(2) it wil give -2sqrt(2) x

OpenStudy (anonymous):

yes... i made a big mistake....

OpenStudy (anonymous):

let's think again

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

could u plz double check the question?

OpenStudy (anonymous):

Ya its same...i checked once again

OpenStudy (anonymous):

Can you lease help me with X school purchased 357 pen and 323 rubber and 289 pencils.The total bill amounts to 4054.5.On the other day school Y purchased 221 pens,247 rubbers and 273 pencils which amounts to 3074.5.Find the price of 100 pens ,100 rubbers and 100 pencils

OpenStudy (anonymous):

Leave that thanks for the effort...try latest one

OpenStudy (anonymous):

Nope i cant solve it o.O

OpenStudy (anonymous):

There must be something wrong with the question.